a body has an initial velocity of 3m/s and has an acceleration of 1m/s^2 direction normal to the of initial velocity. The velocity of the body after 4 sec ?

Grade:12th pass

2 Answers

Basanta Goswami
10 years ago
5m/s
Nirmal Singh.
10 years ago
we will solve this problem in two direction on in is horizontal and one in vertical.....
in horizontal velocity will unchanged after 4 sec i.e Vx = 3 m/s
in vertical motion intial velocity is zero and so final velocity in vertical motion
Vv = uy+at
Vy = 0+1*4 = 4 m/s since acceleration is normal to the initial velocity
so net velocity after 4 sec = v' = (4^2+3^2)^1/2 = 5 m/s
direction theta = tan inverse (vy/vx) = tan inverse (4/3)
Thanks & Regards,
Nirmal Singh
Askiitians faculty